OpenStudy (anonymous):

Rhoel sees ripe mangoes in Jupiter's tree and wants to pick them. he uses a 15 feet long bamboo stick, at an angle of 60 degrees, to poke the mangoes to let it fall. If Rhoel is 5 ft 9 inches tall, how high was the mangoes from the ground?

OpenStudy (anonymous):

\[\sin 60=\frac{opp.}{15}\implies opp=15\sin60=\frac{15\sqrt{3}}{2}\]\[\frac{15\sqrt{3}}{2}+5.75\approx18.74ft\approx18'9"\]